OpenStudy (anonymous):

What are the dimensions of the following rectangle? 3x+2 x A=120 x 3x+2 A. -6 2/3, -18 B. -6 2/3, 6 C. 6 D. 6, 20

OpenStudy (anonymous):

@johnweldon1993 can you help me??

OpenStudy (anonymous):

@formerlyadinosaur can you help me??

OpenStudy (johnweldon1993):

Wait hang on...so does the rectangle look like 

OpenStudy (anonymous):

Yep, that's what it looks like. :-)

OpenStudy (johnweldon1993):

Oh okay.... Alright well first...what do we know about the area of a triangle? \[\large Area = Length \times Width\] correct?

OpenStudy (johnweldon1993):

Area of a rectangle!! **** lol typo there

OpenStudy (anonymous):

Yep, :-)

OpenStudy (johnweldon1993):

Alright....so we know that the length here will be 3x + 2 and the width would be 'x'....so we just plug those into that equation \[\large Area = (3x + 2) \times x\]

OpenStudy (johnweldon1993):

Now we know the total Area would be the 120 they gave....so we can plug that in too...now we have \[\large 120 = (3x + 2) \times x\] good so far?

OpenStudy (anonymous):

Yep, so far. :-)

OpenStudy (johnweldon1993):

Alright now lets simplify the right side there.... \[\large (3x + 2) \times x\] If we distribute that 'x' into both terms...we will have \[\large 3x^2 + 2x\] right?

OpenStudy (anonymous):

That's what I got.

OpenStudy (johnweldon1993):

Okay...so we have \[\large 3x^2 + 2x = 120\] Well this just looks like a quadratic here...which can either be solved by the quadratic formula...or completing the square...which would you prefer?

OpenStudy (anonymous):

The quadratic.

OpenStudy (johnweldon1993):

Alright...so first...lets rewrite our equation as \[\large 3x^2 + 2x - 120 = 0\] okay?

OpenStudy (anonymous):

Yep, :-)

OpenStudy (johnweldon1993):

So now the quadratic formula looks like \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] look familiar?

OpenStudy (anonymous):

Yep it does. :-)

OpenStudy (johnweldon1993):

In that formula...the 'a' 'b' and 'c' represent the coefficients of the equation we have... \[\large \color \red{3}x^2 + \color \red{2}x - \color \red{120}\] so here.... \[\large a = 3\] \[\large b = 2\] \[\large c = -120\] so lets plug those into the quadratic formula

OpenStudy (johnweldon1993):

\[\large \frac{-2 \pm \sqrt{2^2 - 4(3)(-120)}}{2(3)}\] Can you simplify that down for me?

OpenStudy (anonymous):

OK,

OpenStudy (anonymous):

Yes, just a min please.

OpenStudy (johnweldon1993):

Ever figure this out?

OpenStudy (anonymous):

I got \[-25/3\]

OpenStudy (johnweldon1993):

\[\large \frac{-2 \pm \sqrt{2^2 - 4(3)(-120)}}{2(3)}\] \[\large \frac{-2 \pm \sqrt{2^2 + 1440}}{2(3)}\] \[\large \frac{-2 \pm \sqrt{1444}}{2(3)}\] \[\large \frac{-2 \pm 38}{6}\] so we have 2 equations...but 1 solution is negative so we will ignore it... \[\large \frac{36}{6} = 6\] This solved for our 'x'...now we just plug that in fro 'x' in our length and width \[\large length = 3x + 2 \rightarrow 3(6) + 2 \rightarrow 18 + 2 \rightarrow 20\] And \[\large width = x = 6\] So if that is true...we know length = 20 and width = 6

OpenStudy (anonymous):

Thanks a bunch!!! You are really smart If you ever need help with a question I will try my best at it!!! Bye!! ;-)

OpenStudy (johnweldon1993):

No problem :)